OpenStudy (hafeda):

Curve y=ax^2+bx+c passes through (1,2) and is tangent to y=x at the origin find (a,b,c)?? i have the answer i have forgotten how did i solve it, Someone please explain in a little detail. Thanks

OpenStudy (sunnnystrong):

Since the curve passes through the point (1,2), we can start by substituting that into the equation \[2=a(1)^2+b(1)+c\] \[2=a+b+c\] Next, we know that y=x is tangent to the curve at the origin Slope of the line is 1--> \[\frac{ dy }{ dx }=2ax+b\] @ the origin --> 1=2(a)(0)+b 1=b. Evaluate y @ (0,0) \[0=a(0)^2+b(0)+c\] 0=c Find a--> \[2=a(1)^2+1(1)\] a=1 .... @hafeda so what is the equation of the curve?

OpenStudy (hafeda):

Thank u so much, what equation are u asking for ???

OpenStudy (sunnnystrong):

NP Always happy to help :D \[ y=ax^2+bx+c \] .. if you know a,b,c whats the curve?

OpenStudy (hafeda):

ohh y=x^2+x is the curve